Mr. Zaleva flew his private plane from his office to his companies storage facility, bucking a wind of 20-mph head wind the whole way. He flew home the same day with the same wind at his back. The round trip took 10 hours of flying time. If the plane makes 100 mph in still air, how far is the storage facility from his office.
Time= 10 hours
Rate= there= 100-20=80
and back= 100+20=120
Distance= x

Answer by stanbon(48554)

Mr. Zaleva flew his private plane from his office to his companies storage facility, bucking a wind of 20-mph head wind the whole way. He flew home the same day with the same wind at his back. The round trip took 10 hours of flying time. If the plane makes 100 mph in still air, how far is the storage facility from his office.
-------------
Office to storage DATA:
distance = x miles ; rate = 100+20 = 120 mph ; time = d/r = x/120 hrs.
-----------
return trip DATA:
distance = x miles ; rate = 100-20= 80 mph ; time = d/r = x/80 hrs
--------
EQUATION:
time to + time from = 10 hrs
x/120 + x/80 = 10
[80x+120x]=10*80*100
200x = 10*80*100
x=10*40
x=400 miles (distance the storage facility is from his office)
Cheers,
Stan H.